OpenStudy (anonymous):

Alright, so I`m reviewing my old calculus notes and I can do quotient rule and everything, but I`m trying to do first principal stuff again and I`m getting stuck. This is my problem: Determine the Average Rate of Change from x=2 to x=5 given the function: y = x + 1. I know the answer is one using the power rule, but I`m attempting to do this using first principals only. These are my steps: [f(x+h) - f(x)]/h ([(x + h) + 1] - (x + 1) )/h h/h as h -> 0 = 0 .

OpenStudy (amistre64):

again? why???

OpenStudy (amistre64):

Did you forget that the first principles prove the rules and that you never have to look back at it again?

OpenStudy (amistre64):

the rate of change of a line is simply the slope ....

OpenStudy (anonymous):

I know that, but I'm going to redo the course soon due to a lack of 90% the first time. Thus I need to go back to the basics :/

OpenStudy (amistre64):

f(x) = x + 1 f(a+h) - f(a) a+h+1 -a - 1 h ---------- = ------------ = -- = 1 h h h

OpenStudy (amistre64):

since the slope is 1 to begin with

OpenStudy (anonymous):

I thought it was as h-> 0?

OpenStudy (amistre64):

f(x) = 5x + 1 f(a+h) - f(a) 5a+5h+1 -5a - 1 5h ---------- = ---------------- = -- = 5 h h h

OpenStudy (amistre64):

the hs cancel out in the algebra

OpenStudy (anonymous):

Oh i see.

OpenStudy (anonymous):

I was thinking the constant rule >.<. Thank you very much!

OpenStudy (amistre64):

youre welcome